The sum of three numbers in an arithmetic sequence is 9 and their product is 24. Find the three numbers
Answer by Alan3354(69443)

The sum of three numbers in an arithmetic sequence is 9 and their product is 24. Find the three numbers
-----------
9/3 = 3, the middle number
1, 3, 5 doesn't work
--> 2, 3, 4
==================
(n-x) + n + (n+x) = 3n = 9
n = 3
(n-x)*n*(n+x) = 24
(3-x)*3*(3+x) = 24
9 - x^2 = 8
x = 1
--> 2, 3, 4